OpenStudy (anonymous):

Question:

OpenStudy (anonymous):

Solve the equation: \[3^{2x} = 4^{2-x}\]

OpenStudy (anonymous):

do you know logs ? hint : take ln of both sides

OpenStudy (anonymous):

its \[2x \ln 3 = (3-x)\ln4\]

OpenStudy (anonymous):

good.. so you can solve it now dont you ?

OpenStudy (anonymous):

(isolate x)

OpenStudy (anonymous):

\[\frac{ 2x }{ 3-x } = \frac{ \ln4 }{ \ln3 }\]

OpenStudy (anonymous):

you have to solve for x .. so you better isolate x than what you did

OpenStudy (anonymous):

i.e. \[\frac{ 2x }{ 3-x } = 0\]

OpenStudy (anonymous):

\[2xln3=(3-x)\ln4\] \[2xln3+xln4 = 3\ln4\] \[x(2\ln3+\ln4) = 3\ln4\] \[x=3\ln4/(2\ln3+\ln4)\]